A charged particle is moving in a uniform, constan
Answer #1

3.solution

option B is correct

Magnetic fields do no work, so they cannot change the kinetic energy of the particle, hence the speed cannot increase. The velocity (which is a vector) can change*, for the direction may change, but it's magnitude (the speed) cannot change.

* This happens when the particle is moving about in a circle, which actually happens in this situation. In this case, the direction of the particle changes, but the speed is constant

5.solution

The electric force = qE = eE for an electron. The force is upwards as the electron is negative.

The magnetic force = Bqv = Bev for an electron. Using Fleming's left t hand rule and remembering electron flow is opposite to conventional current flow (as electrons are negative), the direction of the force is downwards (so it will cancel the electric force).

When the two forces are equal eE = Bev, so v = E/B.

v = 3000 V/m / 2.0 T

v = 1500 m/s

7.solution

a) The induced EMF is given by the derivative of the magnetic flux x (-1). In this case, the flux is:
Flux = 6*A*B*cos(w*t), where:
A = 0.040 (m^2)
B = 0.20 (T)
w = 2*pi*60 (rad/sec)
So EMF = - (- N*A*B*w)
= 6*0.040*0.2*2*pi*60
= 18.09 sin(wt)
The maximum is 18.09 (V)

8.solution

(a) equal to resistance